Colloquium

A MILK- URP Lecture Mathematics Inquiry Learning at Kutztown Undergraduate Research Presentations

11:00 a.m. Thursday, April 14, 2011

Lytle Hall 228

On the Rudiments of Linear Regression Analysis

Mr. Shawn M. Santo

B.S.-Mathematics & B.S.B.A.-Finance

Kutztown University Class of 2011

Abstract

The term regression analysis was first used by Francis Galton in the late nineteenth century in connection with his work on the relationship between father’s height and son’s height. Today, regression analysis is performed by individuals in both the physical and social sciences. Many of these individuals perform this type of analysis in programs such as Excel, SPSS, Minitab, and SAS without a true understanding of what is actually occurring.

This talk is designed to provide a theoretical focus on the aspects of simple linear regression. More specifically, the coefficients of the linear regression line will be examined. How are these coefficients derived? Do the coefficients provide an accurate estimate of the true population parameters? These are just a few of the questions that will be answered.

Furthermore, the concept of a ”best” linear unbiased estimator will be discussed and its relevance toward simple linear regression.

10:45 a.m. refreshments served

11:00 a.m. talk begins

The Mathematics Inquiry Learning at Kutztown (MILK) Lecture Series is funded by the generous support of the Academy for Inquiry-Based Learning, The Educational Advancement Foundation, and Mr. Harry Lucas, Jr.

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul S. Ache III, Chair